the difference between two numbers is 16. five times the smaller is the same as 8 less than twice the larger. find the numbers
Answer by wuwei96815(245)

Let the first number equal x
Let the second (smaller) number equal x-16

Now, let us write the equation as stated by the word problem:
5(x-16) = 2x-8

Let's solve the equation:
5(x-16) = 2x-8
5x-80 = 2x-8
5x-2x = 80-8
3x = 72
x = 24
x-16 = 8

I hope tht you were able to follow all of that.